The frequency of vibration f of a mass m suspended from o spring of spring constentKis given by arelation of typef=CMxKy,wherec is adimensionally constent. The value ofx and y are

Grade:Select Grade

1 Answers

Arun
6 years ago
Dear Manish
 
Given relation
f = cmxky
where f = Frequency of vibration
       m = Mass suspended from a spring
        k = Spring constant
        c = Dimensionless quantity 
By putting the dimensions of each quantity on both the sides, we get
[T−1 ] = [M]x [MT−2 ] y
 
On comparing the dimensions of quantities on both sides, we get
x + y = 0 and 2y = 1
 
→ y = 1/2
 
and x = −1/2
 
Regards
Arun (askIITians forum expert)